Mathematics 241 – Multivariable Calculus Quiz 6 – October 25, 2013

Your Name:

Let f(x, y) = sin(x) − sin(y).

A) (10) Find all critical points of f in the rectangle [0, 2π] × [0, 2π].

B) (10) Here is the gradient vector field of this f(x, y) plotted on the same rectangle. Classify each critical point as a local maximum, a local minimum, or a saddle point.